OpenStudy (anonymous):

solve the following equation for X (x-16.5)(x-109)=0

OpenStudy (anonymous):

x=16.5,109

OpenStudy (anonymous):

109

OpenStudy (anonymous):

@rachelmcclellan is correct.

OpenStudy (anonymous):

(x-16.5)(x-109)=0 x-16.5=0 x=16.5 x-109=0 x=109

OpenStudy (anonymous):

how about for... x(x+7)+5(x+7)=0

OpenStudy (anonymous):

(x+7)(x+5)=0 x=-7,-5

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

yeaa umm... xsquared+2x=0

OpenStudy (anonymous):

x=0,-2

OpenStudy (anonymous):

x(x+2)=0 x=-2

OpenStudy (anonymous):

how bout.. 4xsquared-8x=0

OpenStudy (anonymous):

x=0,2

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

do u mind helping me with a few more :/ im really bad at this.

OpenStudy (anonymous):

i don't mind at all (:

OpenStudy (anonymous):

4x(x-2)=0 x=2

OpenStudy (anonymous):

I needa solve for X 24x^3-18x^2=0

OpenStudy (anonymous):

x=0, 3/4

OpenStudy (anonymous):

\[6^{2}(4x-3)=0\]

OpenStudy (anonymous):

8x^3-8x^2=0

OpenStudy (anonymous):

6x^2(4x-3)=0 4x-3=0 x=3/4

OpenStudy (anonymous):

I need help with the question above

OpenStudy (anonymous):

whcih one

OpenStudy (anonymous):

the question is 8x^3-8x^2=0

OpenStudy (anonymous):

8x^2(x-1)=0 x-1=0 x=1

OpenStudy (anonymous):

x=0,1

OpenStudy (anonymous):

x(x+2)+3(x+2)=0

OpenStudy (anonymous):

(x+2)(x+3)=0 x=-2,-3

OpenStudy (anonymous):

x=-3,-2

OpenStudy (anonymous):

how bout 2x(x+4)+4(x+4)=0

OpenStudy (anonymous):

(x+4)(2x+4)=0 x=-4,-2

OpenStudy (anonymous):

9(x-10)+x(x-10)=0

OpenStudy (anonymous):

(x-10)(9+x)=0 x=10,-9

OpenStudy (anonymous):

xsquared-11x=0

OpenStudy (anonymous):

x(x-11)=0 x=0,11

OpenStudy (anonymous):

5x(5x+35)-20(5x+35)=0

OpenStudy (anonymous):

(5x+35)(5x-20)=0 x=-7,4

OpenStudy (anonymous):

7(9+x)-7x(9+x)=0

OpenStudy (anonymous):

x=-9,1

OpenStudy (anonymous):

r u done helping me? :/

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ima keep on giving u questions until u tell me to stop lol :)

OpenStudy (anonymous):

(x+5)(x-5)=0

OpenStudy (anonymous):

okay stop